OpenStudy (jmartinez638):

I need help with the evaluation of this limit:

OpenStudy (jmartinez638):

\[\lim_{h \rightarrow 0} \frac{ (x+h)^2\cos(x+h) -x^2\cos(x)}{ h }\]

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{(x+h)^2\left(\sin x\cos h+\sin h\cos x\right)-x^2\cos x}{h}}\) I would start from expanding it a bit

OpenStudy (solomonzelman):

But, this is a lot to write out, what I would do is to go a regular way: \[\lim_{h \rightarrow 0} \frac{ (x+h)^2\cos(x+h) -x^2\cos(x)}{ h }=\frac{d}{dx}\left[x^2\cos x\right]=2x\cos x-x^2\sin x \]

OpenStudy (jhannybean):

How did that all get reduced to \[\frac{d}{dx}(x^2\cos(x))\]

zepdrix (zepdrix):

J, have you learned your derivative shortcut rules at this point? Product rule, chain rule? If not, then we need to do what Solomon first suggested, expanding out all of the mess, we'll group it in a clever way, and relate the pieces to known limits.

zepdrix (zepdrix):

Sorry that was directed at Jmartinez, not Johnnybean hehe

OpenStudy (jhannybean):

Hahaa

OpenStudy (jhannybean):

Oh! can't we simply substitute in 0 for x? :P

zepdrix (zepdrix):

Recall our limit definition for a derivative,\[\large\rm \frac{d}{dx}f(x)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h}\]So we actually have:\[\large\rm \frac{d}{dx}x^2 \cos x=\lim_{h\to0}\frac{ (x+h)^2\cos(x+h) -x^2\cos(x)}{h}\] We clearly can't plug in h=0 right now, it's causing a problem. Plug in x=0? Uhh.. :o

zepdrix (zepdrix):

That gives you \(\large\rm \frac{d}{dx}x^2 \cos x\) at \(\large\rm x=0\). But we want the general formula

zepdrix (zepdrix):

Martinezzzzzz answer the question >.< Have you learned your derivative shortcut rules yet? Power rule, product rule, quotient rule, chain rule. Or are we not up to that point in the course yet? :o It's really important for this problem to know where you are at.

OpenStudy (jhannybean):

Ohhh... I see what he did now, @SolomonZelman reverted back to the original form and took the derivative. I got it!

OpenStudy (jmartinez638):

I have learned those, yes.

OpenStudy (solomonzelman):

I like chain and product better:) All that limit mess is not for me

zepdrix (zepdrix):

Then for this problem, unless the instructions say otherwise, simply relate it back to the function and apply product rule! :)\[\large\rm \lim_{h\to0}\frac{ (x+h)^2\cos(x+h) -x^2\cos(x)}{h}=\frac{d}{dx}x^2 \cos x\]

zepdrix (zepdrix):

Is that confusing? :o

OpenStudy (jmartinez638):

Not really, no xD